# TED: Accelerate Model Training by Internal Generalization

### v2

@@ 3.1 Problem Definition
Given a dataset $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$ containing $n$ training samples, the goal of dataset pruning is to identify a redundant subset $\hat{\mathcal{D}}$ from the original samples, where $\hat{\mathcal{D}}\subset\mathcal{D}$, to accelerate the training process of the model. We can express the optimization parameter through $\mathcal{D}-\hat{\mathcal{D}}$ as $\hat{\theta}_{-\hat{\mathcal{D}}}=\arg\min_{\theta}L(\mathcal{D}-\hat{\mathcal{D}},\theta)$, where $L(\mathcal{D}-\hat{\mathcal{D}},\theta)=\mathbb{E}_{z\in\mathcal{D}-\hat{\mathcal{D}}}loss(z,\theta)$, and $loss$ is the loss function, which is cross-entropy loss for classification tasks.
@@ 3.2 Internal Generalization
Where $\nabla_{\theta_{t-1}}L(\mathcal{D}-\hat{\mathcal{D}},\theta_{t-1})\in\mathbb{R}^{1\times P}$,$\eta$ represents the learning rate. Additionally, the training dynamics are approximated as if they are in continuous time, a method also used in other research [17, 31]. When $n\gg m$, the change in $L(\mathcal{D},\theta_{t})$ at time $t$ is defined, and according to the chain rule, we can obtain:
@@
[…]
@@
$IGD_{-\hat{\mathcal{D}}}=L(\mathcal{D},\hat{\theta}_{-\hat{\mathcal{D}}})-L(\mathcal{D},\hat{\theta})$ (4)
@@
Where $\mathcal{R}(\hat{\theta})$ is the expected loss, and ^$\hat{\mathcal{R}}(\mathcal{D},{\hat{\theta}}_{\mathcal{D}_{k}})$ is the empirical risk of ${\hat{\theta}}_{\mathcal{D}_{k}}$, which is our fitting loss. $\varepsilon$ is a coefficient related to the model size and the size of the retained dataset. When $\hat{\theta}_{D_{k}}$ has the largest upper bound on generalization error, then ^$\hat{\mathcal{R}}(\mathcal{D},{\hat{\theta}}_{{\mathcal{D}}_{k}})$ is also at its maximum. It can be expressed as:
$\mathcal{D}_{k}=\arg\max_{\mathcal{D}_{i}}L(\mathcal{D},\hat{\theta}_{\mathcal{D}_{i}})-L(\mathcal{D},\hat{\theta})$ (7)
@@
$\mathcal{D}_{k}=\arg\min_{\mathcal{D}_{i}}L(\mathcal{D},\hat{\theta}_{-\mathcal{D}_{i}})-L(\mathcal{D},\hat{\theta})$ (8)
@@ 3.3 Dynamic Pruning
$\left|\mathcal{V}_{t}-\mathcal{V}_{T}\right|\leq\left|\nabla_{\theta_{t}}L(\mathcal{D},\theta_{t})\right|C$ (15)

### v3

@@ 3.1 Problem Definition
Given a dataset $\mathcal{D}=\{z_{1},\ldots,z_{n}\}$ containing $n$ training samples, the goal of dataset pruning is to identify a redundant subset $\hat{\mathcal{D}}$ from the original samples, where $\hat{\mathcal{D}}\subset\mathcal{D}$, to accelerate the training process of the model. We can express the optimization parameter through $\mathcal{D}-\hat{\mathcal{D}}$ as $\hat{\theta}_{-\hat{\mathcal{D}}}=\arg\min_{\theta}L(\mathcal{D}-\hat{\mathcal{D}},\theta)$, where $L(\mathcal{D}-\hat{\mathcal{D}},\theta)=\mathbb{E}_{z\in\mathcal{D}-\hat{\mathcal{D}}}loss(z,\theta)$, and lossloss is the loss function, which is cross-entropy loss for classification tasks.
@@ 3.2 Internal Generalization
Where $\nabla_{\theta_{t-1}}L(\mathcal{D}-\hat{\mathcal{D}},\theta_{t-1})\in\mathbb{R}^{1\times P}$$\eta$ represents the learning rate. Additionally, the training dynamics are approximated as if they are in continuous time, a method also used in other research [17, 31]. When $n\gg m$, the change in $L(\mathcal{D},\theta_{t})$ at time $t$ is defined, and according to the chain rule, we can obtain:
@@
$IG=\int_{1}^{T}\frac{dL(\mathcal{D},\theta_{t})}{dt}dt=L(\mathcal{D},\theta_{T})-L(\mathcal{D},\theta_{1})$ (3)
@@
[…]
@@
[…]
@@
[…]
@@ 3.3 Dynamic Pruning
$\left|\mathcal{V}_{t}-\mathcal{V}_{T}\right|\leq\left|\nabla_{\theta_{t}}L(\mathcal{D},\theta_{t})\right|C\$(15)